OpenStudy (anonymous):

Can someone give me examples of a dimension of a cone and a cylinder that have the same volume?

OpenStudy (misty1212):

i don't know much about this, but it cannot be too hard because you can pick any number you like as the volume

OpenStudy (misty1212):

i would make it say \(10\pi\) or something

OpenStudy (misty1212):

then make the dimensions work

OpenStudy (anonymous):

use the equations for volume of a cone and a cylinder\[V_{cone}=\pi r^{2}\frac{ h }{ 3 }\] \[V_{cylinder}=\pi r^{2} h\]

OpenStudy (anonymous):

set them equal to each other and then solve for one of the variables and you can plug in and find the value you want

OpenStudy (mrnood):

you can see that either the height of th ecylinder must be 1/3 that of cone OR r^2 cylinder is 1/3 r^cone OR some combination of the 2 options

OpenStudy (triciaal):

volume of cylinder = 3* volume of cone the easiest way to have the same volume is to have the same radius and the height of the cone 3*height of the cylinder see formula above

OpenStudy (triciaal):

what question do you have about the answers suggested?

OpenStudy (triciaal):

did you try any?

OpenStudy (anonymous):

\[\pi R^2 H=\frac{ 1 }{ 3 }\pi r^2 h\] \[3R^2H=r^2h\] R=radius of cylinder H=height of cylinder. r=radius of cone h=height of cone keeping this equation in mind you can adjust the values yourself.